# Alignment-Sensitive Minimax Rates for Spectral Algorithms with Learned Kernels

### v2

@@ 1 Introduction
Neural networks excel across many applications, yet a complete theoretical understanding of their efficiency remains an open problem. In the infinite-width limit, the Neural Tangent Kernel (NTK) theory approximates training dynamics as kernel regression [23, 2], and it enables the study of generalization by leveraging the classical theory of kernel regression and Reproducing Kernel Hilbert Spaces (RKHS) [8, 48]. However, the NTK theory does not explain why finite-width networks, which adapt their features during training, often outperform traditional methods [19, 16, 25, 37, 43, 36].
A growing line of work directly studies adaptivity, i.e., learning representations or kernel properties during training [4, 27, 31, 10, 47, 51]. Simplified models show that learning eigenvalues (with eigenfunctions fixed) can align the kernel with the signal and improve performance [28, 29]. The common thread is signal-kernel alignment: performance improves when the target’s energy concentrates on leading eigenfunctions [3, 44, 26, 33]. However, classical RKHS theory relies heavily on fixed spectral assumptions (e.g., polynomial eigenvalue decay) and specific signal regularity conditions (e.g., source conditions) [15], which do not apply to learned spectra. To explain the observed advantages of adaptive kernel methods, we need a refined theoretical framework that goes beyond fixed-kernel assumptions.
In this paper, we propose the Effective Span Dimension (ESD), which is a population complexity measure for the analysis of signal-kernel alignment. ESD counts the smallest number of leading eigenfunctions required so that the remaining signal energy matches the estimation variance. Unlike classical measures that ignore the signal, ESD depends on the signal, spectrum, and noise level. Our framework provides new theoretical insights that are absent in classical analyses. In particular, we achieve the following:
@@
    We establish a sharp minimax optimal convergence rate using ESD.
@@
    We explain how gradient-based learning algorithms such as the one in Li and Lin [28] adaptively achieve superior generalization by reducing the ESD.
@@
    We extend our definitions and theory from sequence models to linear regression and kernel regression, which demonstrates the broad applicability of our framework.
Our ESD framework bridges fixed-kernel theory and adaptive learning by quantifying signal-kernel alignment. We hope it will open avenues for deeper understanding of neural networks and novel adaptive algorithms.
Notations. Write $a\lesssim b$ if there exists a constant $C>0$ such that $a\leq Cb$, and write $a\asymp b$ if $a\lesssim b$ and $b\lesssim a$, where the dependence of the constants on other parameters is determined by the context. For $d\in\mathbb{N}_{+}$, let $d]=\{1,2,\ldots,d\}$; for $d=\infty$, let $d]=\mathbb{N}_{+}$. $\mathbf{1}_{\{\cdot\}}$ denotes an indicator function.
@@ 2 Background on Kernel Methods
We therefore seek a refined theoretical framework that explicitly captures signal-kernel alignment and explain the gains achieved by kernel adaptation.
@@
For large $n$, we have $n^{-1}\sum_{i}\psi_{j}(\bm{x}_{i})\psi_{k}(\bm{x}_{i})\approx\mathbb{E}[\psi_{j}(\bm{x})\psi_{k}(\bm{x})]=\mathbf{1}_{\{j=k\}}$, which implies that
@@
This reduction connects RKHS regression to a sequence model where the observations are $z_{j}=\theta^{*}_{j}+\xi_{j}$ and the noise terms $\{\xi_{j}\}$ are uncorrelated with variance $\sigma^{2}:=n^{-1}\sigma_{0}^{2}$. The error in the approximation due to finite $n$ will inflate the estimation variance compared to the sequence model. This approximation error can be controlled if $f^{*}$ is bounded; see Section˜8 for a rigorous treatment.
@@ 3 Effective Span Dimension and Span Profile
where $d\in\{\infty\}\cup\mathbb{N}_{+}$, $\bm{\theta}^{*}=(\theta_{j}^{*})_{j=1}^{d}$ is a sequence of unknown parameters, $\xi_{j}$’s are uncorrelated random variables with mean zero and variance $\sigma^{2}$ (the noise level). For an estimator $\widehat{\bm{\theta}}=(\widehat{\theta}_{j})_{j=1}^{d}$, we consider the loss $\mathcal{L}(\widehat{\bm{\theta}};\bm{\theta}^{*})=\sum_{j=1}^{d}(\widehat{\theta}_{j}-\theta_{j}^{*})^{2}$ and risk $\mathcal{R}(\widehat{\bm{\theta}};\bm{\theta}^{*})=\mathbb{E}\mathcal{L}(\widehat{\bm{\theta}};\bm{\theta}^{*})$. The sequence model captures core estimation phenomena while permitting explicit analysis [11, 24]. In Section˜7, we use whitening to deal with correlated noise and analyze fixed-design linear regression. In Section˜8, we leverage the approximation in Equation˜3 to analyze RKHS regression and random-design linear regression.
@@
Ridge (R): $\psi_{\nu}^{\operatorname{R}}(\lambda)=\frac{1}{1+\lambda/\nu},$ $\widehat{\theta}_{j}^{\operatorname{R},\nu}$ $=\frac{\lambda_{j}}{\lambda_{j}+\nu}\,z_{j}.$ (5)
Gradient Flow (GF): $\psi_{\nu}^{\operatorname{GF}}(\lambda)=e^{-\lambda/\nu},$ $\widehat{\theta}_{j}^{\operatorname{GF},\nu}$ $=\left(1-e^{-\lambda_{j}/\nu}\right)z_{j}.$ (6)
Principal Component (PC): $\psi_{\nu}^{\operatorname{PC}}(\lambda)=\mathbf{1}_{\{\lambda<\nu\}},$ $\widehat{\theta}_{j}^{\operatorname{PC},\nu}$ $=\mathbf{1}_{\{\lambda_{j}\geq\nu\}}\,z_{j}.$ (7)
For spectral estimators, the risk decomposes into squared bias $\sum_{j}\bigl(\psi_{\nu}(\lambda_{j})\bigr)^{2}\,\theta_{j}^{2}$ and variance $\sum_{j}\bigl(1-\psi_{\nu}(\lambda_{j})\bigr)^{2}\,\sigma^{2}$, where $\nu$ controls the bias-variance trade-off. Classical analyses often assume $\bm{\theta}^{*}$ lies in an ellipsoid $\Theta_{\mathbf{a}}=\left\{\theta:\sum_{j}^{\infty}a_{j}^{2}\theta_{j}^{2}\leq C^{2}\right\}$ and derives convergence rates for sequences with $a_{i}\asymp i^{\alpha}$ [24]. Our theoretical framework aims to bypass these assumptions.
@@ 3.1 Effective Span Dimension
Our goal is to develop a measure that captures the interplay between signal structure $\bm{\theta}^{*}$, spectrum $\bm{\lambda}$, and noise variance $\sigma^{2}$. To start, we examine the Principal Component (PC) estimator analytically. PC operates by truncating coordinates with small eigenvalues. Its risk is composed of variance from the retained components and squared bias from those truncated. By trading variance against tail bias, PC admits the optimal truncation point. This motivates our core definition.
@@
$d^{\dagger}=d^{\dagger}(\sigma^{2};\bm{\theta}^{*},\bm{\lambda})=\min\{k\in[d]:\frac{1}{k}\sum_{i=k+1}^{d}\left(\theta_{\pi_{i}}^{*}\right)^{2}\leq\,\sigma^{2}\}.$
Intuitively, the ESD $d^{\dagger}$ is the number of leading coordinates (with leading eigenvalues $\lambda_{i}$) that are most critical for estimation at a given noise level $\sigma^{2}$. It is the truncation point where the squared tail bias of the PC estimator first becomes comparable to (or less than) the estimation variance. The next theorem shows that $d^{\dagger}$ describes the best achievable risk for the PC estimator.
Theorem 3.2 (Optimal PC Estimator Risk).
Let $\widehat{\bm{\theta}}^{\operatorname{PC},\nu}$ be the PC estimator for the sequence model in Equation˜4. Denote by $\mathcal{R}_{*}^{\operatorname{PC}}$ the minimal possible risk over all choices of $\nu$. Let $d^{\dagger}=d^{\dagger}(\sigma^{2};\bm{\theta}^{*},\bm{\lambda})$ be the ESD of $\bm{\theta}^{*}$ w.r.t. the spectrum $\bm{\lambda}$ and the variance $\sigma^{2}$. It holds that
@@
The well-tuned PC estimator is known to be minimax rate optimal under classical assumptions in sequence models that are analogous to the polynomial eigen-decay condition and the source condition in kernel regression (see Propositions 3.11 and 4.23 of Johnstone [24]). In contrast, Theorem˜3.2 suggests that we can instead use $O(d^{\dagger}\sigma^{2})$ to upper bound the minimax estimation error with no reliance on particular spectral decay or source conditions. Moreover, the following theorem confirms that $d^{\dagger}\sigma^{2}$ indeed characterizes the intrinsic difficulty of estimation.
@@
$\mathcal{F}_{K,\bm{\lambda}}^{(n)}=\Bigl\{\bm{\theta}\in\mathbb{R}^{d}:d^{\dagger}(\sigma^{2};\bm{\theta}^{*},\bm{\lambda})\leq K\Bigr\}.$ (8)
@@
Suppose the sample $\bm{Z}$ is drawn from the sequence model in Equation˜4. We have
$\inf_{\widehat{\bm{\theta}}}\sup_{\bm{\theta}^{*}\in\mathcal{F}_{K,\bm{\lambda}}^{(n)}}\mathcal{R}(\widehat{\bm{\theta}},\bm{\theta}^{*})\asymp K\sigma^{2},$
where $\inf$ is taken over any estimator $\widehat{\bm{\theta}}$ based on $\bm{Z}$.
Theorem˜3.3 considers the minimax risk over $\mathcal{F}_{K,\bm{\lambda}}^{(n)}$, a class of distributions whose ESDs are at most $K$. We interpret $K$ as the quota for ESD: the larger $K$, the larger $\mathcal{F}_{K,\bm{\lambda}}^{(n)}$ and thus the higher the minimax risk. By Theorem˜3.2, it is clear that $2K\sigma^{2}$ is an upper bound on the minimax risk, and we only need to establish a matched lower bound.
Theorem˜3.3 highlights the usefulness of ESD: although we motivate its definition using a specific estimator, it quantifies the best possible (worst-case) performance of any estimator. Therefore, the ESD is a fundamental measure for signal-spectrum alignment.
We emphasize that ESD is a population-level complexity measure that depends on the signal and spectrum. It is not a tuning parameter and does not need to be estimated. Its purpose here is explanatory and comparative: it yields a sharp lower bound on the minimax risk and lets us quantify why adaptive learning can outperform fixed-kernel baselines. This mirrors common practice in statistics: sparsity justifies the Lasso, yet practitioners do not estimate the sparsity level before running the method. Estimating ESD from data is an interesting separate problem but orthogonal to our goals in this work.
@@
There are some well-known measures used in the classical analysis of spectral methods. We discuss the differences between ESD and these measures.
@@
In linear regression, Bartlett et al. [5] analyze the minimum-norm interpolator via the effective rank $r_{k}=\frac{\sum_{i>k}\lambda_{\pi_{i}}}{\lambda_{\pi_{k+1}}}$ (using the relationship in Equation˜3). They define the splitting index $k^{*}=\min\{k\geq 0:\,\sigma^{2}r_{k}\geq b\}$ for some constant $b$ and establish risk bounds using $\sigma^{2}k^{*}$. While $k^{*}$ may resemble ESD since both depend on $\bm{\lambda}$ and $\sigma^{2}$, they differ in two important aspects: (i) $k^{*}$ does not involve the signal and thus cannot measure signal-spectrum alignment; and (ii) $k^{*}$ is tailored to the minimum-norm estimator and does not characterize the minimax risk over a class.
Both $d_{\text{eff }}(\nu)$ and $k^{*}$ are signal-agnostic: they depend on the spectrum $\bm{\lambda}$ (and either $\nu$ or $\sigma^{2}$) only, and therefore remain invariant under any change in the alignment between the signal and the kernel’s eigenfunctions. For instance, if adaptive training improves alignment by reordering the eigenfunctions to better align with the signal while preserving the set of eigenvalues, then both $d_{\text{eff }}(\nu)$ and $k^{*}$ are unchanged. In contrast, the ESD $d^{\dagger}\left(\sigma^{2};\bm{\theta}^{*},\bm{\lambda}\right)$ is signal-aware, because it is defined by the bias-variance crossing for the specific $\bm{\theta}^{*}$. As signal-kernel alignment improves, the ESD decreases. Consequently, the ESD can mechanistically explain the generalization benefits of adaptive kernel learning, a phenomenon that signal-agnostic complexity measures like $d_{\text{eff }}(\nu)$ and $k^{*}$ cannot capture.
Examples.
For the following canonical settings, the optimal PC risk satisfies
$\mathcal{R}_{*}^{\mathrm{PC}}\;\asymp\;\begin{cases}\min\!\left\{\sigma^{\frac{2s\beta}{1+s\beta}},\,d\,\sigma^{2}\right\},&\text{(1) }\lambda_{i}=i^{-\beta},\ \sum_{i=1}^{d}\lambda_{i}^{-s}\theta_{i}^{*2}\leq R,\ \beta,s>0,\\[2.0pt] \min\!\left\{\sigma^{\,2-\frac{2}{\alpha}},\,d\,\sigma^{2}\right\},&\text{(2) }\theta_{i}^{*}=i^{-\alpha/2},\ \alpha>1,\ \{\lambda_{i}\}\ \downarrow,\\[2.0pt] \min\!\left\{d\,\sigma^{2},\,\log\!\big(d\sigma^{2}/\log(d\sigma^{2})\big)\right\},&\text{(3) }d<\infty,\ \theta_{i}^{*}=i^{-1/2},\ \{\lambda_{i}\}\ \downarrow,\\[2.0pt] d\,\min\!\left\{d^{-\alpha},\,\sigma^{2}\right\},&\text{(4) }d<\infty,\ 0<\alpha<1,\ \theta_{i}^{*}=i^{-\alpha/2},\ \{\lambda_{i}\}\downarrow,\end{cases}$ (9)
where $\{\lambda_{i}\}\downarrow$ means $\lambda_{i}$ is decreasing. Details and proofs are deferred to Section˜A.3.
In Setting (1), we may take $\sigma^{2}=\sigma_{0}^{2}/n$ in view of Equation˜3, and then the upper bound becomes $\sigma_{0}^{2}\min\left(n^{-\frac{s\beta}{1+s\beta}},d/n\right)$, which matches the well-known optimal rate under the source condition and the polynomial eigen-decay condition in the case when $d=\infty$. When $d<\infty$, there is a phase transition around $d_{0}\asymp n^{\frac{1}{1+s\beta}}$: if $d\lesssim d_{0}$, the upper bound is $d\sigma_{0}^{2}/n$; if $d\gtrsim d_{0}$, the upper bound is the same as if $d=\infty$.
Section˜9.1 illustrates a sparse signal example where the ESD provides a quantitative comparison of two different spectra while the existing measures like $d_{\text{eff }}(\nu)$ and $k^{*}$ do not. These examples suggest that the notion of ESD allows us not only to recover classical results but also to explore new settings where the classical framework is inapplicable.

### v3

@@ 1 Introduction
Many modern learning procedures are adaptive: during training they update a representation, which effectively changes the induced kernel. Such adaptivity can improve generalization when it increases alignment between the learned kernel and the target function [4, 27, 31, 10, 47, 50]. In contrast, classical theory for fixed kernel regression does not capture this effect. For instance, the Neural Tangent Kernel (NTK) theory approximates training dynamics as kernel regression [23, 2], which enables the study of generalization using the classical regression theory in Reproducing Kernel Hilbert Spaces (RKHS) [8, 48]. However, the NTK approximation is intrinsically non adaptive and therefore cannot explain why finite width networks, whose features evolve during training, often outperform traditional methods [19, 16, 25, 37, 43, 36].
Recent simplified models isolate one tractable form of adaptivity: learning the kernel eigenvalues while keeping the eigenfunctions fixed. In these settings, reshaping the spectrum can increase signal kernel alignment and lead to better performance [29, 30]. More broadly, many analyses point to the same mechanism: generalization improves when the target places more energy on the leading kernel eigenfunctions [3, 44, 26, 33]. However, classical RKHS theory is typically developed under fixed spectral assumptions, such as polynomial eigenvalue decay, together with signal regularity conditions, such as source conditions [15]. If the kernel is learned, these assumptions are difficult to justify. To explain the advantages of adaptive kernel methods, we need a refined theoretical framework that remains valid beyond fixed-kernel assumptions.
In this paper, we propose the Effective Span Dimension (ESD) to bridge fixed-kernel theory and adaptive learning. The ESD is a population complexity measure to quantify signal-kernel alignment. Unlike classical measures that ignore the signal, ESD depends on the signal, spectrum, and noise level. Importantly, it is well defined for any kernel, including adaptive kernels whose eigenvalues and eigenfunctions evolve during training. Our framework provides new theoretical insights that are absent in classical analyses. In particular, we achieve the following:
@@
    We establish sharp minimax optimal convergence rates for sequence models using ESD.
@@
    We extend our definitions and theory to linear regression and kernel regression, which demonstrates the broad applicability of our framework.
@@
    We analyze a tractable form of eigenvalue learning to obtain provable ESD reduction. We also empirically illustrate how the ESD framework applies to adaptive kernels with evolving eigenfunctions.
The ESD quantitatively explains the relationship between signal-kernel alignment and generalization. In our framework, the benefit of adaptive learning is exactly a reduction in ESD. Before adaptive training, the kernel can be badly aligned, so the ESD is large and the target lies in a minimax hard family. After training, the kernel is reshaped so that the ESD becomes smaller, which moves the same target into an easier family where efficient spectral methods can generalize well.
Notations.
Write $a\lesssim b$ if there exists a constant $C>0$ such that $a\leq Cb$, and write $a\asymp b$ if $a\lesssim b$ and $b\lesssim a$, where the dependence of the constants on other parameters is determined by the context. For $d\in\mathbb{N}_{+}$, let $d]=\{1,2,\ldots,d\}$; for $d=\infty$, let $d]=\mathbb{N}_{+}$. $\mathbf{1}_{\{\cdot\}}$ denotes an indicator function.
@@ 2 Background on Kernel Methods
Given the above limitation of classical analyses, we therefore seek a refined theoretical framework that explicitly captures signal-kernel alignment and explains the gains achieved by kernel adaptation.
@@
The law of large numbers implies that for large $n$, $n^{-1}\sum_{i}\psi_{j}(\bm{x}_{i})\psi_{k}(\bm{x}_{i})\approx\mathbb{E}[\psi_{j}(\bm{x})\psi_{k}(\bm{x})]=\mathbf{1}_{\{j=k\}}$, which implies that
@@
This reduction connects RKHS regression to a sequence model where the observations are $z_{j}=\theta^{*}_{j}+\xi_{j}$ and the noise terms $\{\xi_{j}\}$ are uncorrelated with variance $\sigma^{2}:=n^{-1}\sigma_{0}^{2}$. The error in the approximation due to finite $n$ will inflate the estimation variance compared to the sequence model. This approximation error can be controlled if $f^{*}$ is bounded; see Section˜9 for a rigorous treatment.
@@ 3 Effective Span Dimension and Span Profile
where $d\in\{\infty\}\cup\mathbb{N}_{+}$, $\bm{\theta}^{*}=(\theta_{j}^{*})_{j=1}^{d}$ is a sequence of unknown parameters, $\xi_{j}$’s are uncorrelated random variables with mean zero and variance $\sigma^{2}$ (the noise level). For an estimator $\widehat{\bm{\theta}}=(\widehat{\theta}_{j})_{j=1}^{d}$, we consider the loss $\mathcal{L}(\widehat{\bm{\theta}};\bm{\theta}^{*})=\sum_{j=1}^{d}(\widehat{\theta}_{j}-\theta_{j}^{*})^{2}$ and risk $\mathcal{R}(\widehat{\bm{\theta}};\bm{\theta}^{*})=\mathbb{E}\mathcal{L}(\widehat{\bm{\theta}};\bm{\theta}^{*})$. The sequence model captures core estimation phenomena while permitting explicit analysis [11, 24]. In Section˜8, we use whitening to deal with correlated noise and analyze fixed-design linear regression. In Section˜9, we leverage the approximation in Equation˜3 to analyze RKHS regression and random-design linear regression.
@@
$\text{Ridge (R):}\quad\psi_{\nu}^{\operatorname{R}}(\lambda)=\frac{1}{1+\lambda/\nu},\widehat{\theta}_{j}^{\operatorname{R},\nu}=\frac{\lambda_{j}}{\lambda_{j}+\nu}\,z_{j}.$ (5)
$\text{Gradient Flow (GF):}\quad\psi_{\nu}^{\operatorname{GF}}(\lambda)=e^{-\lambda/\nu},$
$\qquad\qquad\qquad\quad\qquad\widehat{\theta}_{j}^{\operatorname{GF},\nu}=\left(1-e^{-\lambda_{j}/\nu}\right)z_{j}.$ (6)
$\text{Principal Component (PC):}\quad\psi_{\nu}^{\operatorname{PC}}(\lambda)=\mathbf{1}_{\{\lambda<\nu\}},$
$\qquad\qquad\qquad\qquad\qquad\qquad\widehat{\theta}_{j}^{\operatorname{PC},\nu}=\mathbf{1}_{\{\lambda_{j}\geq\nu\}}\,z_{j}.$ (7)
For spectral estimators, the risk decomposes into squared bias $\sum_{j}\bigl(\psi_{\nu}(\lambda_{j})\bigr)^{2}\,\theta_{j}^{2}$ and variance $\sum_{j}\bigl(1-\psi_{\nu}(\lambda_{j})\bigr)^{2}\,\sigma^{2}$, where $\nu$ controls the bias-variance trade-off. Classical analyses often assume $\bm{\theta}^{*}$ lies in an ellipsoid $\Theta_{\mathbf{a}}=\left\{\theta:\sum_{j}^{\infty}a_{j}^{2}\theta_{j}^{2}\leq C^{2}\right\}$ and derive convergence rates for sequences with $a_{i}\asymp i^{\alpha}$ [24]. Our theoretical framework aims to bypass these assumptions.
@@ 3.1 Effective Span Dimension
Our goal is to quantify how estimation difficulty depends jointly on the signal $\bm{\theta}^{*}$, the spectrum $\bm{\lambda}$, and the noise variance $\sigma^{2}$. We first introduce the effective span dimension (ESD) as an intrinsic, population-level quantity determined by $(\bm{\theta}^{*},\bm{\lambda},\sigma^{2})$ through the ordering of the spectrum.
@@
$d^{\dagger}(\sigma^{2};\bm{\theta}^{*},\bm{\lambda})=\min\{k\in[d]:\frac{1}{k}\sum_{i=k+1}^{d}\left(\theta_{\pi_{i}}^{*}\right)^{2}\leq\,\sigma^{2}\}.$
When $\lambda_{j}$ is decreasing, the summation in the above definition is simply the sum of the squared tail of $\bm{\theta}^{*}$. Using the indices $\pi_{i}$, the definition applies to general ordering.
Intuitively, the ESD $d^{\dagger}$ is the smallest number $k$ of leading eigencoordinates (ordered by $\bm{\lambda}$) needed so that the remaining signal energy falls below $k\sigma^{2}$. It is a complexity index of the pair (signal, spectrum) at noise level $\sigma^{2}$, and it exists regardless of which algorithm is used for estimation.
We next show that $d^{\dagger}$ describes the best achievable risk for the PC estimator, which serves as an achievability argument for the rates that will be shown minimax optimal in Theorem˜3.3.
Theorem 3.2.
Let $\widehat{\bm{\theta}}^{\operatorname{PC},\nu}$ be the PC estimator in Equation˜7 for the sequence model in Equation˜4. Denote by $\mathcal{R}_{*}^{\operatorname{PC}}$ the minimal possible risk over all choices of $\nu$. Let $d^{\dagger}=d^{\dagger}(\sigma^{2};\bm{\theta}^{*},\bm{\lambda})$ be the ESD of $\bm{\theta}^{*}$ w.r.t. the spectrum $\bm{\lambda}$ and the variance $\sigma^{2}$. It holds that
@@
In classical theories on sequence models, the well-tuned PC estimator is known to be minimax rate optimal under regularity assumptions that are analogous to the polynomial eigen-decay condition and the source condition in kernel regression (see Propositions 3.11 and 4.23 of Johnstone [24]). In view of Theorem˜3.2, those minimax rates should be at the order of $O(d^{\dagger}\sigma^{2})$. Indeed, we can use the ESD to characterize the intrinsic difficulty of estimation, which is the goal of the following theorem.
@@
$\mathcal{F}_{K,\bm{\lambda}}^{(n)}=\Bigl\{\bm{\theta}\in\mathbb{R}^{d}:d^{\dagger}(\sigma^{2};\bm{\theta},\bm{\lambda})\leq K\Bigr\}.$ (8)
@@
Suppose the sample $\bm{Z}$ is drawn from the sequence model in Equation˜4 and consider any estimator $\widehat{\bm{\theta}}$ based on $\bm{Z}$. We have
$\inf_{\widehat{\bm{\theta}}}\sup_{\bm{\theta}^{*}\in\mathcal{F}_{K,\bm{\lambda}}^{(n)}}\mathcal{R}(\widehat{\bm{\theta}},\bm{\theta}^{*})\asymp K\sigma^{2}.$
The class $\mathcal{F}_{K,\bm{\lambda}}^{(n)}$ consists of distributions whose ESDs are at most $K$. We interpret $K$ as the quota for ESD: the larger $K$, the larger $\mathcal{F}_{K,\bm{\lambda}}^{(n)}$. Theorem˜3.3 states that, over $\mathcal{F}_{K,\bm{\lambda}}^{(n)}$, the minimax risk grows linearly with $K$. Theorem˜3.3 confirms that the ESD quantifies the best possible performance of any estimator over the class $\mathcal{F}_{K,\bm{\lambda}}^{(n)}$.
Example 3.4.
For any given spectrum and $\bm{\theta}^{*}$, we can find the smallest $K^{*}$ such that $\bm{\theta}^{*}\in\mathcal{F}_{K^{*},\bm{\lambda}}^{(n)}$, over which the minimax risk is $\sigma^{2}K^{*}$. Below are several examples:
  • (1)
    If $\lambda_{i}=i^{-\beta}$ and $\sum_{i=1}^{d}\lambda_{i}^{-s}\theta_{i}^{*2}\leq R$ for some $\beta,s>0$,
    $\sigma^{2}K^{*}\;\asymp\;\min\!\left\{\sigma^{\frac{2s\beta}{1+s\beta}},\,d\sigma^{2}\right\}.$
  • (2)
    If $\theta_{i}^{*}=i^{-\alpha/2}$ for some $\alpha>1$ and $\{\lambda_{i}\}_{i=1}^{d}$ is decreasing, then
    $\sigma^{2}K^{*}\;\asymp\;\min\!\left\{\sigma^{2-\frac{2}{\alpha}},\,d\sigma^{2}\right\}.$
  • (3)
    If $d<\infty$, $\theta_{i}^{*}=i^{-1/2}$, and $\{\lambda_{i}\}_{i=1}^{d}$ is decreasing, then
    $\sigma^{2}K^{*}\;\asymp\;\min\!\left\{d\sigma^{2},\,\log\!\big(d\sigma^{2}/\log(d\sigma^{2})\big)\right\}.$
  • (4)
    If $d<\infty$, $0<\alpha<1$, $\theta_{i}^{*}=i^{-\alpha/2}$, and $\{\lambda_{i}\}_{i=1}^{d}$ is decreasing, then
    $\sigma^{2}K^{*}\;\asymp\;d\min\!\left\{d^{-\alpha},\,\sigma^{2}\right\}.$
Details and proofs are deferred to Section˜A.3.
In Setting (1), we may take $\sigma^{2}=\sigma_{0}^{2}/n$ to align with the approximation developed in Equation˜3, and then the upper bound becomes $\sigma_{0}^{2}\min\left(n^{-\frac{s\beta}{1+s\beta}},d/n\right)$, which matches the well-known optimal rate under the source condition and the polynomial eigen-decay condition in the case when $d=\infty$. When $d<\infty$, there is a phase transition around $d_{0}:=n^{\frac{1}{1+s\beta}}$: if $d\lesssim d_{0}$, the upper bound is $d\sigma_{0}^{2}/n$; if $d\gtrsim d_{0}$, the upper bound is the same as if $d=\infty$.
These examples suggest that the notion of ESD allows us not only to recover classical results but also to explore new settings where the classical framework is inapplicable. ∎
ESD is a population complexity measure for theoretical analysis rather than a training input. Comparing the ESD of the same signal across kernels indicates which kernel can generalize better for that signal. Section˜10.1 illustrates a sparse signal example where the ESD provides a quantitative comparison of two different spectra while the existing measures like $d_{\text{eff }}(\nu)$ and $k^{*}$ do not.
Discussion on estimator-guided measures.
The ESD equals the truncation point for the PC estimator where the squared tail bias first becomes comparable to (or less than) the estimation variance. It may seem that one can define an alternative complexity index by calibrating the bias and variance of any particular estimator. There are two reasons why we focus on the ESD.
First, Theorem˜3.3 clarifies that any complexity measure that characterizes minimax optimal rates for the classes we consider above must agree with ESD up to constant factors. Therefore, ESD is the canonical measure for alignment-sensitive minimax rates. We use ESD because it admits a clean analytic characterization and enables the study of the dynamic signal-kernel alignment in adaptive learning.
Second, such an estimator-guided index may reflect defects of the chosen procedure rather than intrinsic difficulty. A concrete example is ridge regression: due to the saturation effect, a ridge-guided index calibrated to ridge’s bias–variance balance will fail to describe the optimal estimation risk even in classical settings with source conditions. We rigorously formalize this point in Section˜5.
@@
We discuss the differences between ESD and some well-known measures used in the classical analysis of spectral methods.
@@
In linear regression, Bartlett et al. [6] analyze the minimum-norm interpolator via the effective rank $r_{k}=\frac{\sum_{i>k}\lambda_{\pi_{i}}}{\lambda_{\pi_{k+1}}}$ (using the relationship in Equation˜3). They define the index $k^{*}=\min\{k\geq 0:\,\sigma^{2}r_{k}\geq b\}$ for some constant $b$ and establish risk bounds using $\sigma^{2}k^{*}$. $k^{*}$ differs from ESD in two important aspects: (i) $k^{*}$ does not involve the signal and thus cannot measure signal-spectrum alignment; and (ii) $k^{*}$ is tailored to the minimum-norm estimator and does not characterize the minimax risk over a class.
Both $d_{\text{eff }}(\nu)$ and $k^{*}$ are signal-agnostic: they depend on the spectrum $\bm{\lambda}$ (and either $\nu$ or $\sigma^{2}$) only, and therefore remain invariant under any change in the alignment between the signal and the kernel’s eigenfunctions. For instance, if adaptive training improves alignment by reordering the eigenfunctions to better align with the signal while preserving the set of eigenvalues, then both $d_{\text{eff }}(\nu)$ and $k^{*}$ are unchanged. In contrast, the ESD $d^{\dagger}\left(\sigma^{2};\bm{\theta}^{*},\bm{\lambda}\right)$ is signal-aware, because it is defined by the bias-variance crossing for the specific $\bm{\theta}^{*}$. As signal-kernel alignment improves, the ESD decreases. Consequently, the ESD can mechanistically explain the generalization benefits of adaptive kernel learning, while $d_{\text{eff }}(\nu)$ and $k^{*}$ cannot.